In the morning we have a residual Pie-Shape from the LBTI night.

10:30 AdSec recovers. Start of activities with cal. unit, we calibrate the FLAO with the on-axis source.

REC: 20140512_175558: FLAO REC wiht bin 2, TT only, KL_v16

We restore the Yellow reconstructor on ARGOS. FLAO slopes repositioned in the REC to be after # 1200.

We reverse the ARGOS REC to have a positive gain instead of negative (Tommaso reverses all the recs recorded so far).

REC: 150modes_yellow_flaoTTbin2: FLAO + ARGOS yellow reconstructor reversed

PatCam Yellow fails so we put in the Blue pupil.

REC: 150modes_blue_flaoTTbin2: FLAO + ARGOS blue reconstructor reversed

The FLAO+ARGOS loop diverges so we reload just the ARGOS blue REC (still reversed). Also in this way the loop diverges.

12:00 while the ARGOS blue loop is divergin we get a Pie-Shape.
Topic revision: r1 - 12 May 2014, MarcoBonaglia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback